Carbohydrates = 4.1 Calories/gram (includes sugars and starches; insoluble fiber cannot be digested by humans, and does not provide calories for humans)Protein = 4.1 Calories/gramLipids = 9.5 Calories/gram

A). A one ounce serving of potato chips has 10 g of fat and 13 g of carbohydrates, mostly in the form of starches, but one g of carbohydrates is insoluble fiber. How many calories are in this serving?

b). How many calories are in a 5 ounce potato chip serving(as described in #a)? For a moderately active 55 year-old man, what is the percentage of calories relative to the total daily requirement of a 5 ounce serving of potato chips?
Answer #1

ANSWER.

A) Calculation of calories in the serving.

       Carbohydrate present in 1 ounce of potato chips

                                                             = 13 gm.(1gm is insoluble fiber).

          1 gm of carbohydrate = 4.1 calories.

      Insoluble fiber not provide calories.

         So

           Calories present in 13 gm of carbohydrate = 12 * 4.1

                                                                                = 49.2 calories.

    Fat or lipids present in 1 ounce of potato chips = 10 gm.

              1 gm of lipid = 9.5 calorie.

         Calories present in 10 gm of lipid = 9.5 * 10

                                                                = 95 calories.

Total amount of calories in 1 ounce of potato chips = 49.2 + 95

                                                                                    = 144.2 calories.

B) Calories in a 5 ounce of potato chips.

         Total amount of calories in 1 ounce of potato chips = 144.2 calories.

           So

            Calories present in 5 ounce of potato chips = 144.2 * 5

                                                                                  = 721 calories.

   * Percentage of calories relative to the total daily requirement of 5 ounces

       of potato chips.

           calories present in 5 ounce of carbohydrate = 49.2 * 5

                                                                                    = 246 calories.

              Percentage of calories of carbohydrate

                           in 5 ounces of potato chips = 246 / 721 * 100

                                                                             = 34 %

          Calories present in 5 ounce of fat = 95 * 5

                                                               = 475 calories.

           Percentage of calories of fat

                           in 5 ounce of potato chips = 475 / 721 * 100

                                                                   = 65.8 %